log378= 2.5775
-------------------
5In(r) + 3In(t) - 4In(s)
= lnr^5 + lnt^3 - lns^4
= ln[r^5*t^3/s^4]
----------------------------

In(x) +In(x-2) = In(x+2) + In(x-3)
g(x) = xto the 2nd power + 3 and h(x) = x+ 3
______
2 Find (h o g)(x)

Answer by stanbon(57361)

log378= 2.5775
-------------------
5In(r) + 3In(t) - 4In(s)
= lnr^5 + lnt^3 - lns^4
= ln[r^5*t^3/s^4]
----------------------------

In(x) +In(x-2) = In(x+2) + In(x-3)
= ln[x(x-2)] = ln[(x+2)(x-3)]
x(x-2) = (x+2)(x-3)
x^2 - 2x = x^2 -x -6
x = 6
--------------------------------
g(x) = x^2+3 and h(x) = (x+ 3)/2

Find (h o g)(x) = h[x^2+3] = [(x^2+3)+3]/2 = [x^2+6]/2
===============================
Cheers,
Stan H.